OpenStudy (anonymous):

Molly works at an electronics store. Her last customer purchased a DVD player for $69.80 and five DVDs for $61.12, including taxes. If the customer paid her with two $100 bills, how much change should she give him? A. $69.08 B. $60.28 C. $130.20 D. $138.88

OpenStudy (anonymous):

@jim_thompson5910 @tkhunny

OpenStudy (anonymous):

first do 69.80+61.12

OpenStudy (tkhunny):

What's preventing you from adding the two sale prices and subtracting the sum from $200?

OpenStudy (anonymous):

130.92

OpenStudy (anonymous):

a.

OpenStudy (tkhunny):

Move on to the next one.

OpenStudy (anonymous):

Silas took 13 bags of glass to the recycling center. He still has 11 bags of plastic to take to the recycling center. Which equation could be used to find x, the total number of bags of glass and plastic Silas will take to the recycling center? A. 13 - x = 11 B. x - 11 = -13 C. x - 13 = 11 D. x + 13 = 11

OpenStudy (anonymous):

a is correct because the customer would give her $200.00. then you would subtract 130.92 from 200.00 and get 69.08

OpenStudy (anonymous):

oh that was for the first question not the second lol

OpenStudy (tkhunny):

1) Put different questions on different threads. 2) Never post a question with no work shown (or personal ideas presented) 3) We started with 13 and ended up with 11. Anything look like it says that?

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

c and a

OpenStudy (anonymous):

im not sure if it is c or a

OpenStudy (anonymous):

oh! it is c because x represents the amount of bags of glass he started with. then he took 13 to the recycling bin so that is (-13) and he ended up with 11 he still needed to take. so x-13=11

OpenStudy (anonymous):

i hope that helps

OpenStudy (tkhunny):

That's it. 'A' was tempting until I realized there were two types of bags.

OpenStudy (anonymous):

Thank you!!!!:)